# On the Adversarial Robustness of LASSO Based Feature Selection

### v1

@@ III Markov Blanket and Feature Selection
Proposition 1 $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$ iif $D_{wass}[p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}),p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})]=0$. where $D_{wass}[\mathbf{p},\mathbf{q}]$ denotes the Wasserstein distance between distribution $\mathbf{p}$ and $\mathbf{q}$.
@@
Proof: Suppose the elements of the distance matrix $\mathbf{D}$ in Problem(3) are positive except the diagonal elements and let $\mathbf{p}(Y)=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}),\mathbf{q}(Y)=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. The solution of the optimization problem(3) is zero if and only if the joint distribution matrix $Q_{i,j}=0$ for all $i\neq j$. Then we have $\mathbf{Q}\cdot\mathbf{1}=\mathbf{Q}^{\top}\mathbf{1}$, so $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. Also, if $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. It is easy to see that the optimal solution of problem(3) is zero.
@@
Proposition 1 tells us that the Wasserstein distance can reveal the independence property in the probabilistic graphic model.
@@
In the paper [7], they tried to find a Markov Blanket for each feature $X_{i}$. However, finding the Markov Blanket for a feature is too hard in practice. Proposition 2 tells us a more practical way to find whether $X_{i}$ belongs to the Markov Blanket.
@@
Proposition 2 Let $\mathbf{X}_{\mathcal{G}_{i}}$ denotes a subset of $\mathbf{X}\setminus X_{i}$ and contains all the features which $X_{i}$ directed to. If for every setting of $Y=y,\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}$ we have $p(Y=y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y=y|\mathbf{X}_{\mathcal{G}_{i}})$. Then, $X_{i}$ does not belong to the Markov Blanket.
@@
Proof: Since $X_{i}$ and $Y$ are independent given $\mathbf{X_{\mathcal{G}_{i}}}$ and $\mathbf{X}_{\mathcal{G}_{i}}$ contains all the nodes which $X_{i}$ direct to, if $X_{i}$ belongs to the Markov Blanket, there should be a node which both $Y$ and $X_{i}$ directed to, which contradicts the assumption that $Y$ and $X_{i}$ are independent given $\mathbf{X}_{\mathcal{G}_{i}}$.
@@
As [7] has pointed out, even if we can find the appropriate feature set $\mathbf{X}_{\mathcal{G}_{i}}$, in practice, the distance of the two conditional probabilities,$p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})$ and $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$ , will hardly be zero. This is due to the noise or the limited samples. To test whether $p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}})$, we need compare every configuration of $\mathbf{X}_{\mathcal{G}_{i}},X_{i}$. To make it more convenient and take the weights of the $\mathbf{X}_{\mathcal{G}_{i}},X_{i}$ into consideration, we use the averaged distance to measure their discrepancy. The averaged distance is defined as:
$\delta(X_{i})=\mathbb{E}_{\mathbf{X}_{\mathcal{G}_{i}},X_{i}}D_{wass}[p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i}),p(Y|\mathbf{X}_{\mathcal{G}_{i}})]$ (4)
@@
Here is our algorithm:
  1. 1.
@@
    Compute the absolute value of correlation $\rho_{i,j}=|\frac{Cov(X_{i},X_{j})}{D(X_{i})D(X_{j})}|$ for each pair of feature $X_{i}$ and $X_{j}$. Let $\mathbf{\bar{X}}=\mathbf{X}$.
  2. 2.
@@
    For every $X_{i}\in\mathbf{\bar{X}}$, find the features with $L$ largest value of $\rho_{i,j}$ and define them as $\mathbf{X}_{\mathcal{G}_{i}}$.
  3. 3.
    Compute $\delta(X_{i})$ for each $X_{i}\in\mathbf{\bar{X}}$.
  4. 4.
@@
    Find $X_{i}$ which has the smallest $\delta(X_{i})$, delete $X_{i}$ from $\mathbf{\bar{X}}$, and let $\mathbf{\bar{X}}=\mathbf{\bar{X}}\setminus X_{i}$.
  5. 5.
    Return to (1) until some stopping criteria are satisfied.
IV Feature Selection with Noisy Labels
$\underset{|\theta|=K}{\text{argmin}}\,\mathbb{E}_{\mathbf{X}}D_{wass}[p(\tilde{Y}|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X}_{\theta})]$ (5)
where $D_{wass}[\mathbf{p},\mathbf{q}]$ is Wasserstein distance between $\mathbf{p}$ and $\mathbf{q}$. Here, unlike most of the analysis, we do not assume a specific label noisy model[33] in our analysis. We still use the set $\theta\subseteq\{1,2,\cdots,M\}$ to denote the selected feature set. Fig.(3) shows the relationships between the conditional distributions before and after feature selection in the original clean data set as well as in the observed noisy labeled data set.
Before further analysis, we need define some quantities to simplify the notation of our analysis. First, $\epsilon_{1}$ refers to the expected Wasserstein distance between the original conditional distribution and the noisy labeled conditional distribution.
$\epsilon_{1}=\mathbb{E}_{\mathbf{X}}D_{wass}[p(Y|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X})]$
Second, $d_{1}(\theta)$ refers to the expected Wasserstein distance between the conditional distribution before feature selection and after feature selection in the original data samples.
$d_{1}(\theta)=\mathbb{E}_{\mathbf{X}}D_{wass}[p(Y|\mathbf{X}),\,p(Y|\mathbf{X}_{\theta})]$
It tells us if we select the feature set $\theta$ and how far we deviate from the original conditional distribution. Similarly, $d_{2}(\theta)$ means the distance in the noisy labeled data set before feature selection and after feature selection:
$d_{2}(\theta)=\mathbb{E}_{\mathbf{X}}D_{wass}[p(\tilde{Y}|\mathbf{X}),\,p(\tilde{Y}|\mathbf{X}_{\theta})]$
In addition, $\epsilon_{2}$ is defined as the expected Wasserstein distance between the conditional distribution in the original clean data sample and the conditional distribution in the noisy labeled data set after feature selection:
$\epsilon_{2}=\mathbb{E}_{\mathbf{X}_{\theta}}D_{wass}[p(Y|\mathbf{X}_{\theta}),\,p(\tilde{Y}|\mathbf{X}_{\theta})]$
$\theta_{1}^{*}=\underset{|\theta|=K}{\text{argmin}}\quad d_{1}(\theta)$ (6)
$\theta_{2}^{*}=\underset{|\theta|=K}{\text{argmin}}\quad d_{2}(\theta)$ (7)
$\theta_{1}^{*}$ indicates the optimally selected feature set using the original data and $\theta_{2}^{*}$ indicates the optimally selected feature set using the noisy labeled data. Although $\theta_{2}^{*}$ is the best feature set we can get based on the noisy labeled data set, to give a quantitative evaluation of the optimally selected features based on the noisy labeled data set, we use the quantity
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})$ (8)
Conclusion 1: For the same selected feature set $\theta$ in the original data and in the noisy labeled data, $\epsilon_{2}\leq\epsilon_{1}$.
TABLE I: The left optimization problem calculates the distance from the original data to the noisy labeled data; the right optimization problem calculates the distance from the original data to the noisy labeled data after feature selection.
Proof: To compute $\epsilon_{1}$ which involves solving the left optimization problem of Table(LABEL:table1-two-opts) for each configuration of the random vector $\mathbf{X}=\mathbf{x}$ and to compute $\epsilon_{2}$ involves solving the right optimization problem of Table(I). We denote $\mathbf{p}_{Y|\mathbf{x}}=[p(Y=c_{1}|\mathbf{X}=\mathbf{x}),p(Y=c_{2}|\mathbf{X}=\mathbf{x}),\cdots,p(Y=c_{n}|\mathbf{X}=\mathbf{x})]^{\top}$ and the similar vector notation for $\mathbf{p}_{Y|\mathbf{x}_{\theta}},\mathbf{p}_{\tilde{Y}|\mathbf{x}}$, and $\mathbf{p}_{\tilde{Y}|\mathbf{x}_{\theta}}$.
Suppose we have gotten the optimal solution for the left problem as $\mathbf{Q}(\mathbf{x})$, where the optimal solution is denoted as a function of $\mathbf{x}$. So, the optimal value of the left problem is $\sum_{i,j}Q(\mathbf{x})_{i,j}D_{i,j}$. By taking expectation over the random variable $\mathbf{X}$, we get the expression for the value of $\epsilon_{1}$ to be $\epsilon_{1}=\int p(\mathbf{x})\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j}\,\mathrm{d}\mathbf{x}$.
Another matrix $\mathbf{\tilde{Q}}(\mathbf{x}_{\theta})$ with entries defined as
$\tilde{Q}_{i,j}(\mathbf{x}_{\theta})=\frac{\int p(\mathbf{x})Q_{i,j}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
First, we use Bayesian rule to calculate $p(Y=c_{i}|\mathbf{x}_{\theta})$:
$p(Y=c_{i}|\mathbf{x}_{\theta})=\frac{\int p(Y=c_{i}|\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$ (9)
According to the first constraint in the left problem, we have:
$p(Y=c_{i}|\mathbf{x})=\sum_{j}Q_{i,j}(\mathbf{x})$
Plug this equation into Eq.(9) and we get
$p(Y=c_{i}|\mathbf{x}_{\theta})$ $=\frac{\int\sum_{j}Q_{i,j}(\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\frac{\int Q_{i,j}(\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\tilde{Q}_{i,j}(\mathbf{x}_{\theta})$
Accordingly, following the same routine above, we get
$p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})=\frac{\int p(\tilde{Y}=c_{i}|\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$p(\tilde{Y}=c_{i}|\mathbf{x}_{\theta})$ $=\frac{\int p(\mathbf{x})\sum_{j}Q_{j,i}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\frac{\int p(\mathbf{x})Q_{j,i}(\mathbf{x})\,\mathrm{d}\mathbf{x}_{\bar{\theta}}}{p(\mathbf{x}_{\theta})}$
$=\sum_{j}\tilde{Q}_{j,i}(\mathbf{x})$
$\epsilon_{2}$ $=\mathbb{E}_{\mathbf{x}_{\theta}}D_{wass}[p(Y|\mathbf{X}_{\theta})\,,p(\tilde{Y}|\mathbf{X}_{\theta})]$
$\leq\int\sum_{i,j}\tilde{Q}_{i,j}(\mathbf{x}_{\theta})D_{i,j}p(\mathbf{x}_{\theta})\,\mathrm{d}\mathbf{x}_{\theta}$
$=\int\sum_{i,j}\frac{\int Q_{i,j}(\mathbf{x})p(\mathbf{x})\,\mathrm{d}\mathbf{x_{\bar{\theta}}}}{p(\mathbf{x}_{\theta})}p(\mathbf{x}_{\theta})D_{i,j}\,\mathrm{d}\mathbf{x}_{\theta}$
$=\int\sum_{i,j}Q_{i,j}(\mathbf{x})D_{i,j}p(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\epsilon_{1}$
It demonstrates that $\epsilon_{2}$ is always less than or equal to $\epsilon_{1}$. By constraining selecting the same feature set under the original data and under the noisy labeled data, the value of $\epsilon_{2}$ is constrained to be no larger than $\epsilon_{1}$.
$-\epsilon_{1}-\epsilon_{2}\leq d_{1}(\theta_{1}^{*})-d_{2}(\theta_{1}^{*})\leq\epsilon_{1}+\epsilon_{2}$ (10)
$-\epsilon_{1}-\epsilon_{2}\leq d_{1}(\theta_{2}^{*})-d_{2}(\theta_{2}^{*})\leq\epsilon_{1}+\epsilon_{2}$ (11)
$d_{1}(\theta_{2}^{*})\geq d_{1}(\theta_{1}^{*})$ (12)
$d_{2}(\theta_{1}^{*})\geq d_{2}(\theta_{2}^{*})$ (13)
Combine inequalities (10)(11)(12)(13) and we have
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})+d_{2}(\theta_{1}^{*})-d_{2}(\theta_{2}^{*})\leq 2\epsilon_{1}+2\epsilon_{2}$ (14)
$d_{1}(\theta_{2}^{*})-d_{1}(\theta_{1}^{*})\leq 4\epsilon_{1}$ (15)
This inequality shows that the optimal feature set $\theta_{2}^{*}$ under noisy labeled data is $4\epsilon_{1}$ suboptimal in the original clean data.
To compute $\epsilon_{1}$, we need to solve the optimization problem in the left of Table(I) for every configuration of $\mathbf{X}=\mathbf{x}$. Because this optimization problem does not have a closed form solution in general cases, we can not get a closed form solution for $\epsilon_{1}$. Also, the relationship between $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ is not specified, which makes this problem even harder to solve. Instead of directly solving this problem, we can find some meaningful bound for $\epsilon_{1}$.
$p(\tilde{Y}|\mathbf{X})=\sum_{Y}p(\tilde{Y}|Y)p(Y|\mathbf{X})$ (16)
where $p(\tilde{Y}|\mathbf{Y})$ denotes the transition probability which quantifies the the probability that the true label is $Y$ while observed is $\tilde{Y}$. Under this model, it is still hard to get closed form solution of $\epsilon_{1}$. However, we can construct a joint distribution matrix and get it’s upper bound. Let the elements of joint distribution matrix be $Q_{i,j}(\mathbf{x})=p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})$ and $p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})=p(\tilde{Y}=c_{j}|Y=c_{i})p(Y=c_{i}|\mathbf{x})$. We can verify that this matrix is a feasible solution of the left problem of Table(I). So, we have
$=\int p(\mathbf{x})\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})D_{i,j}\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$ (17)
The NNAR model assumes that the observed label depends on both the true label and its features. So, the observed class conditional distribution given the features is:
$p(\tilde{Y}|\mathbf{X})=\sum_{Y}p(\tilde{Y}|Y,\mathbf{X})p(Y|\mathbf{X})$ (18)
Compared with Eq.(16), it only differs in the form of the transition probability. Construct a feasible joint probability matrix as $Q_{i,j}(\mathbf{x})=p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})$ and $p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{x})=p(\tilde{Y}=c_{j}|\mathbf{Y}=c_{i},\mathbf{x})p(Y=c_{i}|\mathbf{x})$. Under this noise model, we can also find an upper bound of $\epsilon_{1}$:
$\epsilon_{1}$ $\leq\int p(\mathbf{x})\sum_{i,j}p(Y=c_{i},Y=c_{j}|\mathbf{x}),\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$ (19)
Generally, let $p(Y,\tilde{Y}|\mathbf{X})$ denotes the joint conditional distribution of $Y$ and $\tilde{Y}$ given feature $\mathbf{X}$, where $Y$ is the original label and $\tilde{Y}$ is the noisy label and it satisfies the constraints of the optimization problem, that is, its marginal conditional distributions are $p(Y|\mathbf{X})$ and $p(\tilde{Y}|\mathbf{X})$ respectively. Then we get an upper bound for optimal value of the optimization problem:
$\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})D_{i,j}$
By taking the expected value regarding $\mathbf{X}$, we get the upper bounded solution of $\epsilon_{1}$:
$\epsilon_{1}$ $\leq\int p(\mathbf{X}=\mathbf{x})\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})D_{i,j}\,\mathrm{d}\mathbf{x}$
$=\sum_{i,j}D_{i,j}\int p(\mathbf{X}=\mathbf{x})p(Y=c_{i},\tilde{Y}=c_{j}|\mathbf{X}=\mathbf{x})\,\mathrm{d}\mathbf{x}$
$=\sum_{i,j}p(Y=c_{i},\tilde{Y}=c_{j})D_{i,j}$ (20)
V The algorithm complexity analysis

### v3

@@ II-C Markov Blanket
$p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}})$ iif $W[p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i}),p(Y|\mathbf{X}_{\mathcal{G}_{i}})]=0$ for every $\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}$, where $\mathcal{G}_{i}$ is a subset of the features, $i\notin\mathcal{G}_{i}$, and $W[\mathbf{p},\mathbf{q}]$ denotes the Wasserstein distance between distribution $\mathbf{p}$ and $\mathbf{q}$.
Proof.
@@
Suppose the elements of the distance matrix $\mathbf{D}$ in Problem(3) are positive except the diagonal elements and let $\mathbf{p}(Y)=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}),\mathbf{q}(Y)=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. The solution of the optimization problem(3) is zero if and only if the joint distribution matrix $Q_{i,j}=0$ for all $i\neq j$. Then we have $\mathbf{Q}\cdot\mathbf{1}=\mathbf{Q}^{\top}\mathbf{1}$, so $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. Also, if $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$. It is easy to see that the optimal solution of problem(3) is zero. ∎
@@
In Proposition 1, $p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}})$ implies $p(Y,X_{i}|X_{\mathcal{G}_{i}})=p(Y|X_{\mathcal{G}_{i}})\cdot p(X_{i}|\mathbf{X}_{\mathcal{G}_{i}})$. It tells us that the Wasserstein distance can reveal the independence property in the probabilistic graphic model.
@@
In paper [7], they tried to find a Markov Blanket for each feature $X_{i}$. However, finding the Markov Blanket for a feature is too hard in practice. Proposition 2 tells us a more practical way to find whether $X_{i}$ belongs to the Markov Blanket.
Proposition 2.
@@
Let $\mathbf{X}_{\mathcal{G}_{i}}$ denotes a subset of $\mathbf{X}\setminus X_{i}$ and contains all the features which $X_{i}$ directed to. If for every setting of $Y=y,\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i}$ we have $p(Y=y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})=p(Y=y|\mathbf{X}_{\mathcal{G}_{i}})$. Then, $X_{i}$ does not belong to the Markov Blanket.
Proof.
@@
Since $X_{i}$ and $Y$ are independent given $\mathbf{X_{\mathcal{G}_{i}}}$ and $\mathbf{X}_{\mathcal{G}_{i}}$ contains all the nodes which $X_{i}$ direct to, if $X_{i}$ belongs to the Markov Blanket, there should be a node which both $Y$ and $X_{i}$ directed to, which contradicts the assumption that $Y$ and $X_{i}$ are independent given $\mathbf{X}_{\mathcal{G}_{i}}$. ∎
@@
As [7] has pointed out, even if we can find the appropriate feature set $\mathbf{X}_{\mathcal{G}_{i}}$, in practice, the distance of the two conditional probabilities, $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}},X_{i}=x_{i})$ and $p(Y|\mathbf{X}_{\mathcal{G}_{i}}=\mathbf{x}_{\mathcal{G}_{i}})$ , will hardly be zero. This is due to the noise or the limited samples. To test whether $p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i})=p(Y|\mathbf{X}_{\mathcal{G}_{i}})$, we need compare every configuration of $\mathbf{X}_{\mathcal{G}_{i}}$ and $X_{i}$. To make it more convenient and consider the weights of $\mathbf{X}_{\mathcal{G}_{i}}$ and $X_{i}$ into consideration, we use the averaged distance to measure their discrepancy. The averaged distance is defined as:
$\delta(X_{i})=\mathbb{E}_{\mathbf{X}_{\mathcal{G}_{i}},X_{i}}W[p(Y|\mathbf{X}_{\mathcal{G}_{i}},X_{i}),p(Y|\mathbf{X}_{\mathcal{G}_{i}})].$ (4)
@@
II-D Feature Selection Algorithm
With help of Proposition 1 and Proposition 2, here we give our algorithm.
First, compute the absolute value of correlation $\rho_{i,j}=|\frac{Cov(X_{i},X_{j})}{std(X_{i})std(X_{j})}|$ for each pair of feature $X_{i}$ and $X_{j}$. Those values are used in the following step to estimate the distance between two feature nodes in a probabilistic graphic model. Let $\mathbf{\bar{X}}=\mathbf{X}$ to denote the current existing features.
Second, For every $X_{i}\in\mathbf{\bar{X}}$, find the features with $L$ largest value of $\rho_{i,j}$ and define them as $\mathbf{X}_{\mathcal{G}_{i}}$. In this step, we are trying to find all the nodes which $X_{i}$ directed to. According to the ”folk-theorem”[7], the closer of two nodes, the influence tend to be stronger. By choosing $L$ largest correlated nodes, we believe it will contain all the node $X_{i}$ directed to.
Third, Compute $\delta(X_{i})$ for each $X_{i}\in\mathbf{\bar{X}}$. According to Proposition 1 and Proposition 2, if $\delta(X_{i})$ is zero, $X_{i}$ does not belong to the Markov Blanket.
Fourth, Find $X_{i}$ which has the smallest $\delta(X_{i})$, delete $X_{i}$ from $\mathbf{\bar{X}}$, and let $\mathbf{\bar{X}}=\mathbf{\bar{X}}\setminus X_{i}$. Because of the noise or limited sample effect, $\delta(X_{i})$ will hardly be zero even if it is not contained in the Markov Blanket. So, for each run, we choose the one with smallest $\delta$ value and delete it from the remaining feature set.
We summarize our algorithm in Table II-D.
TABLE I: summary of our algorithm
{tabu}
@@ II-D Feature Selection Algorithm
to 0.45 X[l] 1. Compute the absolute value of correlation $\rho_{i,j}=|\frac{Cov(X_{i},X_{j})}{std(X_{i})std(X_{j})}|$ for each pair of feature $X_{i}$ and $X_{j}$. Let $\mathbf{\bar{X}}=\mathbf{X}$.
@@
2. For every $X_{i}\in\mathbf{\bar{X}}$, find the features with $L$ largest value of $\rho_{i,j}$ and define them as $\mathbf{X}_{\mathcal{G}_{i}}$.
3. Compute $\delta(X_{i})$ for each $X_{i}\in\mathbf{\bar{X}}$.
@@
4. Find $X_{i}$ which has the smallest $\delta(X_{i})$, delete $X_{i}$ from $\mathbf{\bar{X}}$, and let $\mathbf{\bar{X}}=\mathbf{\bar{X}}\setminus X_{i}$.
5. Return to (1) until the size of the $\bar{\mathbf{X}}$ equal to $K$.
II-E Complexity Analysis